The volume of a rectangular package is 2304 cubic inches. The lenght of the package is 3 times its width, and the height is 1.5 times its width. Find the dimensions of the package.
Answer by checkley77(12844)

L=3W
H=1.5W
L*H*W=2,304
3W*1.5W*W=2,304
4.5W^3=2,304
W^3=2,304/4.5
W^3=512
W=CUBERT512
W=8 IS THE WIDTH OF PACKAGE.
L=3*8=24 IS THE LENGTH OF PACKAGE.
H=1.5*8=12 IS THE HEIGHT OF PACKAGE.
PROOF:
8*24*12=2,304
2,304=2,304